# Perturbation-Restrained Sequential Model Editing

### v1

@@ A.2 Preliminary Lemmas & Theorems
which establishes (37). Also $\hat{\kappa}\leq\frac{\kappa}{\gamma}$, and the inequality (38) follows from (25).
@@ A.3 Core Theorem
To bound the first term in (45), note that
@@
For the second term in (45) we have
@@
The bound (41) follows on combining (42)–(47).
Readers can refer to this work [51] for more details of perturbation analysis.
@@ B.1 Baseline Editing Methods
    MEND [43] [^6] https://github.com/eric-mitchell/mend: it learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form.
@@
    ROME [41] [^7] https://github.com/kmeng01/rome: it first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module.
@@
    MEMIT [42] [^8] https://github.com/kmeng01/memit: it extended ROME to edit a large set of facts and updated a set of MLP layers to update knowledge.
The ability of these methods was assessed based on EasyEdit [^9] https://github.com/zjunlp/EasyEdit [58], an easy-to-use knowledge editing framework which integrates the released codes and hyperparameters from previous methods.
@@ B.2 Editing Datasets and Evaluation Metrics
Table 2 shows the examples of two datasets.
Table 2: The editing datasets of both CounterFact and ZsRE.
@@
ZsRE Which was the record label for New Faces, New Sounds?
Besides, following previous works [41, 43, 42], the editing performance metrics for the CounterFact dataset and ZsRE dataset are efficacy, generalization and locality, but there are some computational differences. In the main paper, the metrics of editing performance are calculated for the CounterFact dataset. For the ZsRE dataset, here are the details:
@@ C.1 Results of General Abilities
Figure 6 and 7 show the downstream task performance of edited models with GPT-2 XL and LLaMA-3 (8B) on CounterFact dataset. Figure 8 shows the downstream task performance of edited models with LLaMA-2 (7B) on ZsRE dataset. Due to limitations of computing resources, experiments were conducted using only LLaMA-2 (7B) on the ZsRE dataset. We will supplement experiments with other LLMs in the future.
@@
Figure 8: The downstream task performance (%) of models edited by three editing methods with LLaMA-2 (7B) on the ZsRE dataset.
@@ C.2 Results of Editing Performance
Figure 9 and 10 shows the editing performance of edited models with GPT-2 XL and LLaMA-3 (8B) on CounterFact dataset. Figure 11 shows the editing performance of edited models with LLaMA-2 (7B) on ZsRE dataset.
Figure 9: The editing performance (%) of three editing methods with GPT-2 XL on CounterFact dataset.
@@
Figure 11: The editing performance (%) of editing methods with LLaMA-2 (7B) on the ZsRE dataset.
@@ C.3 Results of another function for PRUNE
In the main paper, $\operatorname{log}$ function is used in $F$ in PRUNE to restrain $\hat{\sigma}_{i}$. Here we use the $\operatorname{linear}$ function, which could be represented as: $F(\hat{\sigma}_{i})=\frac{1}{\beta}*\hat{\sigma}_{i}+\frac{\beta-1}{\beta}*max\{\sigma_{i}\}$. Here $\beta>1$ was a hyperparameter and was set as 2 in this section. Figure 12 and 13 respectively show some downstream task performance and editing performance with $\operatorname{linear}$ function on CounterFact dataset.
Compared with Figure 6 and 9, we observed that although the $\operatorname{linear}$ function in PRUNE played a role in preserving general abilities and maintaining editing performance, its effectiveness was noticeably inferior to that of the $\operatorname{log}$ function when the number of edits was large.
@@ Appendix D Broader Impacts
By considering both the positive and negative impacts and implementing appropriate mitigation strategies, this work can contribute to the responsible and ethical advancement of model editing technologies.

### v3

@@ A.2 Preliminary Lemmas & Theorems
which establishes equation 37. Also $\hat{\kappa}\leq\frac{\kappa}{\gamma}$, and the inequality equation 38 follows from equation 25.
@@ A.3 Core Theorem
To bound the first term in equation 45, note that
@@
For the second term in equation 45 we have
@@
The bound equation 41 follows on combining equation 42–equation 47.
Readers can refer to this work (Stewart & Sun, 1990) for more details of perturbation analysis.
@@ B.1 Baseline Editing Methods
    MEND (Mitchell et al., 2022a) [^6] https://github.com/eric-mitchell/mend: it learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form.
@@
    ROME (Meng et al., 2022) [^7] https://github.com/kmeng01/rome: it first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module.
@@
    MEMIT (Meng et al., 2023) [^8] https://github.com/kmeng01/memit: it extended ROME to edit a large set of facts and updated a set of MLP layers to update knowledge.
The ability of these methods were assessed based on EasyEdit [^9] https://github.com/zjunlp/EasyEdit (Wang et al., 2023), an easy-to-use knowledge editing framework which integrates the released codes and hyperparameters from previous methods.
@@ B.2 Editing Datasets and Evaluation Metrics
Table 3 shows the examples of two factual datasets (ZsRE) (Levy et al., 2017) and CounterFact (Meng et al., 2022). Figure 6 shows an example of ConceptEdit dataset, which is cited from Wang et al. (2024). More details can refer to the original paper of these datasets.
Table 3: The editing datasets of both ZsRE and CounterFact.
@@
ZsRE Which was the record label for New Faces, New Sounds?
@@
Figure 6: An example of ConceptEdit dataset
Besides, following previous works (Meng et al., 2022; Mitchell et al., 2022a; Meng et al., 2023), the editing performance metrics for the ZsRE and CounterFact datasets are efficacy, generalization and locality, but there are some computational differences. In the main paper, the metrics of editing performance are used for the ZsRE dataset.
For the CounterFact dataset, here are the details:
@@ C.1 Results of General Abilities
Figure 7,  8 and 9 show the downstream task performance of edited models with GPT-2 XL, LLaMA-2 (7B) and LLaMA-3 (8B) on CounterFact dataset. Due to limitations of computing resources, experiments were conducted using only LLaMA-2 (7B) on the ZsRE dataset. We will supplement experiments with other LLMs in the future.
@@
Figure 8: The downstream task performance (%) of models edited by three editing methods with LLaMA-2 (7B) on the CounterFact dataset.
@@ C.2 Results of Editing Performance
Figure 10,  11 and 12 shows the editing performance of edited models with GPT-2 XL, LLaMA-2 (7B) and LLaMA-3 (8B) on CounterFact dataset.
Figure 10: The editing performance (%) of three editing methods with GPT-2 XL on CounterFact.
Figure 11: The editing performance (%) of three editing methods with LLaMA-2 (7B) on the CounterFact dataset.
@@ C.3 Results of another function for PRUNE
In the main paper, $\operatorname{log}$ function is used in $F$ in PRUNE to restrain $\hat{\sigma}_{i}$. Here we use the $\operatorname{linear}$ function, which could be represented as: $F(\hat{\sigma}_{i})=\frac{1}{\beta}*\hat{\sigma}_{i}+\frac{\beta-1}{\beta}*max\{\sigma_{i}\}$. Here $\beta>1$ was a hyperparameter and was set as 2 in this section. Figure 13 and 14 respectively show some downstream task performance and editing performance with $\operatorname{linear}$ function on CounterFact dataset.
Compared with Figure 7 and 10, we observed that although the $\operatorname{linear}$ function in PRUNE played a role in preserving general abilities and maintaining editing performance, its effectiveness was noticeably inferior to that of the $\operatorname{log}$ function when the number of edits was large.
@@ C.4 Condition number with PRUNE
C.5 The correlation between condition number and general abilities
Figure 16 simultaneously shows the condition number and general abilities of three editing methods without PRUNE in the sequential editing process. From these experiments, we observed that a dramatic increase in the condition number is often accompanied by a rapid decline in general abilities.
Figure 16: The condition number and downstream task performance of three editing methods with LLaMA-2 (7B) on the CounterFact dataset. Since MEMIT and MEND have multiple parameters to be edited, we randomly selected one of them to calculate the condition number.
C.6 Enhanced PRUNE
We find applying multiple operation of PRUNE for longer-term performances will perform better than only once. The results of applying PRUNE once have been shown in our paper. Here, we make a comparison. For convenience, we list some results of using ROME to edit GPT 2-XL on the CounterFact dataset.
Table 6: comparison of applying once operation of PRUNE and applying multiple multiple operation.
Mode General Abilities Editing Performance
Method Edits Reasoning Summa Open-QA NLI Efficacy General Locality
ROME+PRUNE (once) 100 30.12 12.12 22.63 49.83 99 92 74.10
500 27.26 10.78 17.86 45.34 56.80 46.05 72.32
ROME+PRUNE (multiple) 100 30.17 11.38 22.99 48.83 100 98 76.1
500 32.68 11.52 22.99 51.17 84.6 78.7 75.1